How to find 32^2 /5  .pls provide the  exact steps .(ans is 2)

Grade:11

1 Answers

ROSHAN MUJEEB
3 years ago
Dear Student
Please check the right answer is 4. 32 is 2^5
Therefore 2^5*2/5 = 2^2=4
Hope you got it